# LEPTODERMIC CORRECTIONS TO THE TOV EQUATIONS AND NUCLEAR ASTROPHYSICS WITHIN THE EFFECTIVE SURFACE APPROXIMATION

### v1

@@ II Tolman-Oppengeimer-Volkoff approach
Notice that the solution, Eq. (11), coincides, up to a constant, with that presented in Tolman’s book at $\zeta=B_{\rm S}/A_{\rm S}=1/(2A_{\rm S})$ (see Refs. RT87 ; MS24 and Eq. (9) for $R_{\rm S}$). The cosmological constant also is assumed to be zero, and the units for the dimensionless pressure are used, $\mathcal{E}_{0}\propto R^{-2}_{\rm S}$, Eq. (9). Thus, one obtains the relation of these parameters of the Schwarzschild metric, Eq. (II), to the initial condition $P=P_{0}$ at $r=0$ through Eq. (12) for $\zeta$. Notice that according to Eq. (11) for the pressure $P$, one has $P=0$ at the boundary $r=R$ because of Eq. (10) for the constants $A_{\rm S}$ and $B_{\rm S}$. Finally, our result for the pressure, Eq. (11), coincides identically with that of Ref. RT87 .
Figure 2 shows the contour plots for the pressure $P(r)$ [in units of $P(r=0)$], as function of $r$ (km) and parameter $A_{\rm S}$, Eq. (10). It is more convenient to use the finite dimensionless values of $A_{\rm S}$, $0\leq A_{\rm S}\leq 1.5$, instead of dimensional $R_{\rm S}$ for the fixed effective NS radius $R=10$ km, $0\leq r\leq R$. For a very rough estimate neglecting the relatively small gravitational mass defect, $\mathcal{E}_{0}\sim c^{2}\rho^{(0)}$, where the central mass density $\rho^{(0)}\approx M/(4\pi R^{3}/3)\sim M_{\odot}/(4\pi R^{3}/3)$, ($M\approx M_{V}$) and $M_{\odot}=1.989*10^{30}$ kg is the Sun mass, one finds significantly larger Schwarzschild radius parameter $R_{\rm S}$ than the effective radius $R$, $R_{\rm S}\approx 15.6$ km for $M=1.4M_{\odot}$, as example. By definition, $R_{\rm S}$ is always larger than (or equal to) the effective NS radius $R$, $R\leq R_{\rm S}$, while the gravitational radius $r_{\rm g}$ should be smaller than $R$, $r_{\rm g}=2GM/c^{2}\leq R$ LLv2 . These conditions are fulfilled in our calculations for the values $R_{\rm S}\approx 15.6$ km, and $r_{\rm g}\approx 4.1$ km in the case of the effective radius $R=10$ km. With these evaluations, one has the constant $A_{\rm S}\approx 1.15$. The maximum value $A_{\rm S}\approx 1.5$ corresponds to zero pressure $P(r)$.
@@ III Modified TOV approach
We now derive the leptodermic corrections to the TOV equations due to the gradient terms of the energy density $\mathcal{E}$, in line of Ref. MS24 . The details of these derivations are shown in Appendix A.
@@ III.1 General points
The key quantity of our derivations is the energy density $\mathcal{E}\left(\rho\right)$ strmagden ; BM15 ; MS24 ,
@@
The volume element, $\hbox{d}\mathcal{V}$, for a spherical system is determined by the Schwarzschild metric, $\hbox{d}\mathcal{V}=\hbox{exp}(\lambda/2)\hbox{d}{\bf r}$, where $\hbox{d}{\bf r}=r^{2}\hbox{sin}\theta~{}\hbox{d}\theta\hbox{d}\varphi$. In Eq. (13), $\mathcal{A}(\rho)$ is the volume non-gradient part, Eq. (9),
@@
where $K$ is the non-gravitational part of the incompressibility modulus, e.g., due to the Skyrme nuclear interaction (for nuclear matter $K\sim 200$ MeV), $\Phi(\rho)$ is the statistically averaged (macroscopic) gravitational potential; see Eq. (8) and discussions around it in Ref. MS24 . For simplicity, we will consider the gradient term of the inter-particle interaction with a main constant $\mathcal{C}$ of Eq. (12) for a smooth coefficient $\mathcal{B}(\rho)$ in front of the gradient-density square term.
@@
Therefore, when we can neglect external forces, there is no linear terms in expansion of $\mathcal{A}$, Eq. (15), over powers of the difference $\rho-\overline{\rho}$ near the saturation value $\overline{\rho}$. Near the saturation value, $\rho\rightarrow\overline{\rho}$, one has $\mathcal{W}=\mathcal{A}/\rho$. Expanding the potential $\Phi(\rho)$ near the saturation density $\overline{\rho}$ up to second order, and taking into account Eq. (17), one indeed finds that there is no linear term, namely, $\partial\Phi/\partial\rho=0$ at $\rho=\overline{\rho}$, i.e., $\Phi=\Phi_{0}+(1/2)\Phi_{2}(\rho-\overline{\rho})^{2}$ with some constants $\Phi_{0}$ and $\Phi_{2}$. Combining different terms in Eq. (15), one writes
$\mathcal{A}=--b^{(G)}_{V}\rho+\varepsilon_{G}(\rho),$ (18)
where $b^{(G)}_{V}=b_{V}-m\Phi_{0}$, $m$ is the particle mass, $\Phi_{0}=\Phi(\overline{\rho})$. The particle separation energy, $b^{(G)}_{V}$, and energy density component, $\varepsilon_{G}(\rho)$ (see Eq. (16) for $\varepsilon(\rho)$) are modified both by the gravitational field $\Phi(\rho)$,
@@
$\frac{\delta\mathcal{E}}{\delta\rho}\equiv\frac{\partial\mathcal{A}}{\partial\rho}-2\mathcal{C}\Delta\rho=\mu~{}.$ (22)
@@
The solution of the Lagrange equation (22) at leading order of the leptodermic expansion over the parameter $a/R$, $\rho(r)=\overline{\rho}y((r-R)/a)$, as function of the radial coordinate $r$, is shown in Fig. 1 in the main case (i), $\mathcal{B}=\mathcal{C}=const$, for typical parameters, the thickness of the NS diffused layer (crust) $a=1$ km and the effective radius $R=10$ km, in units of the saturation value $\overline{\rho}$; see Refs. ST04 ; HPY07 ; MS24 . We compare the two dimensionless solutions for the particle density $\rho(r)/\overline{\rho}$. One of them (solid line) is related to the vdW-Skyrme interaction constant $\mathcal{C}$ “1” for a dense liquid drop; see Eq. (32) for the universal dimensionless particle number density of a dense liquid drop, $y=\rho/\overline{\rho}$, as function of $x=(r-R)/a$. Another limit case corresponds to the symmetric Wilets solution (ii) wilets ; MS24 for zero constant $\mathcal{C}$ and no spin-orbit term but with taking into account the only kinetic gradient term in a gas system; see Eq. (12) for $\mathcal{B}$. These two curves are very different in the surface layer, especially outside of the system far away from the NS, and almost the same inside of the NS. As seen from Fig. 1 and expected, the range of the exponential decrease of $\rho(r)$ becomes much smaller for the vdW-Skyrme liquid-drop case (i) than for the Wilets gas case (ii). But the surface effects on the local density $\rho(r)$ in both cases are still very important to get a continuous analytical behavior through the ES. The smallness factor, proportional to $a/R$, with respect to the volume contribution, appears because of the integration over the normal-to-ES (radial in this example) variable $x=(r-R)/a$ in this ES region. Thus, for the calculations of the surface corrections to the NS total energy $E$, Eq. (14), and particle number $N$, Eq. (21), it would seem possible to assume that they are not (very) important for so small relative diffuseness thickness $a/R=0.1$. However, as shown in Ref. MS24 , the influence of the NS surface becomes really important for the local quantities as the particle density $\rho(r)$, Fig. 1, the EoS, and therefore, for the pressure $P(r)$ calculating through Eqs. (13) and (17),
@@
Moreover, the gravitational mass defect due to the radial curvature of the nonlinear Schwarzschild metric in calculations of the NS mass leads to its dramatic nonmonotonic behavior of $M=M(R)$; see Ref. MS24 .
@@ III.2 Surface corrections to the TOV equations
Indeed, $\mathcal{E}_{1}$ enters the solutions of Eq. (2) in terms of the integral over the radial coordinate $r$. Our derivations at leading order over leptodermic parameter $a/R$ are similar to those for the equation for the density $\rho$ and its solutions in Ref. MS24 .
Starting from the same equations (2), as presented in the previous section (see Ref. RT87 ), we split all their solutions into two components, the leading component and leptodermic corrections,
@@
As shown in Appendix A, using approximately the property of the energy density component $\mathcal{E}_{1}\propto(\partial\rho/\partial r)^{2}$, which is concentrated near the ES, for the correction $\lambda_{1}$, at first order in the leptodermic expansion, one finally obtains from the second equation of Eq. (2),
@@
Substituting the asymptote of the solutions, $\rho\sim\hbox{exp}(-\xi/a)$, for larger distances $\xi/a\gg 1$ (see Ref. MS24 ), one has an exponential disappearance of the integral $\mathcal{D}_{1}$, Eq. (29),
@@
For the interior region, one finds also the exponentially decreasing asymptotes of $\mathcal{D}_{1}$, Eq. (29),
@@
where $y=\rho/\overline{\rho}$ (see Ref. MS24 ), $C_{y}$ is an integration constant. To remove singularity at the origin, we shall assign the value $-4/15$ for the constant $C_{y}$. Notice that in the derivations for $\lambda_{0}$, Eq.(27) [Eq. (96.7) in Tolman’s book RT87 ] similar corrections were neglected by the same arguments. The solution $\lambda_{1}(r)$, Eq. (28), is zero outside of the system because of zero $\mathcal{E}_{0}$, Eq. (6), and $P_{0}$, Eq. (11). Therefore, $\lambda_{1}(r)$ is concentrating near the NS surface, as it should be true (see also Appendix A).
Substituting the derivative ${\rm d}\nu/{\rm d}r$ from the first equation of Eq. (2) to its last equation, for a given energy density $\mathcal{E}$, one obtains the modified TOV equation for the pressure $P$,
@@
where $\lambda=\lambda_{0}+\lambda_{1}$; see Eqs. (27) and (28). This is valid approximately up to high, e.g. second order in the leptodermic expansion. Let us use now Eqs. (24) for $\mathcal{E}$ and (26) for $P$. Using also Eq. (A1) for $P_{0}$ at zero order of the leptodermic parameter $a/R$, one can derive the linearized equation for $P_{1}$ beyond the analytical solutions (27) and (11) of the standard TOV equations for $\lambda_{0}$ and $P_{0}$, respectively.
@@
where $\tilde{r}=r/R_{\rm S}$, as in Appendix A. In these equations, $P_{0}(r)$ is the analytical solution of the TOV equations at zero order of the leptodermic parameter for the pressure, Eq. (11); see Refs. RT87 ; MS24 . Using the same method of the arbitrary constant variations, one obtains the solution in terms of the quadratures (see Appendix A),
@@ III.3 Leading solutions for the pressure
Substituting the function $\beta(r)$, Eq. (35), into the integral for the pressure correction $P_{1}$, Eqs. (37) and (38), one finds
@@
where $\zeta=1/(3\sqrt{1-\tilde{R}^{2}})$. Using then Eq. (36) for $\beta_{1}(r)$ in Eq. (38) for $C_{1}^{(P)}$, one obtains the sum of the two terms,
@@
The integrand of one of them is proportional to $\mathcal{E}_{1}$, Eq. (25), while the second one is proportional to $\lambda_{1}$, Eq. (28). The first order corrections, $\mathcal{E}_{1}$ and $\lambda_{1}$, are determined by Eqs. (25) and (28), respectively [see also the approximate Eq. (A6)]. Both these functions exponentially decrease for $r$ outside of a relatively small range $a$ of the NS crust at small leptodermic parameter, $a/R$. Therefore, the first integral in Eq. (III.3) can be taken approximately at leading order, $a/R$, in the leptodermic expansion as for the derivation of $\lambda_{1}$. Notice that the second integral in Eq. (III.3), with the integrand $\propto\lambda_{1}(r)$, over radial coordinate $r$ can be neglected as the second order of $(a/R)^{2}$ in the leptodermic expansion. Therefore, it does not contribute at first order over $a/R$. Indeed, this integral is reduced approximately to $\int{\rm d}r\mathcal{Y}(y)$,
@@
×$\times\int{\rm d}r\mathcal{Y}(y).$ (41)
Notice that by the same reason, the $\lambda_{1}$ surface correction to the mass defect owing to the curved space can be neglected in calculations of the NS mass $M$ (Ref. MS24 ) at the first order over $a/R$. Transferring the last integral over $r$ to the variable $\rho$, ${\rm d}r={\rm d}\rho({\rm d}r/{\rm d}\rho)$, and using Eqs. (24), and (19) for the inverse derivative of $\rho$ over $r$ (${\rm d}r={\rm d}\xi$), one finds
@@
The integral on the right hand side is finite because its integrand has no singularity in the limit $y\rightarrow 1$ by the L’hospital theorem, $\mathcal{Y}/(1-y)\rightarrow$ $\mathcal{Y}^{\prime}(y)/(-1)\rightarrow(1-y)\rightarrow 0~{}$. There is also no singularity at small $y$, $\int{\rm d}y/\sqrt{y}=2\sqrt{y}$.
@@
where $\mathcal{D}_{1}$ is given by Eq. (29) with the same asymptotic properties (30) and (31) outside of the narrow NS crust, and $\tilde{R}=R/R_{\rm S}$. With Eq. (39) for the integral exponent taken at $r=R$ and the coefficient $C^{(P)}_{1}$, Eq. (43), one arrives at the final analytical approximation for the surface pressure correction $P_{1}$,
$P_{1}=-\frac{3R^{3/2}}{4R_{\rm S}^{5/2}\sqrt{1-R^{2}/R^{2}_{\rm S}}}~{}\mathcal{D}_{1}~{};$ (44)
see Eq. (37). Here, $R$ is the effective radius, $R_{\rm S}$ is the Schwartzschild radius, Eq. (9), related to the initial value of the energy density, $\mathcal{E}_{0}$, and $\mathcal{D}_{1}$ is given above by Eq. (29) with the internal (31) and external (30) asymptotes. Therefore, these leptodermic corrections for $C_{1}^{(P)}$, and hence, for $P_{1}$, (as $\propto\mathcal{D}_{1}$), are exponentially decreasing functions of radial coordinate $r$, $r\rightarrow\infty$ outside [see Eq. (30)] and to zero inside of the system [Eqs. (31) and (III.2)]. In particular, inside of the system, Eq. (31) for $\mathcal{D}_{1}$, one arrives at
$P_{1}=\frac{a}{R}~{}\frac{\overline{\rho}K_{G}}{12}~{}\frac{R^{5/2}}{R^{5/2}_{\rm S}\sqrt{1-R^{2}/R^{2}_{\rm S}}}~{}\mathcal{Y}(y)~{}.$ (45)
Thus, as expected, the leading leptodermic correction $P_{1}$ to the zero order pressure $P_{0}$ is peaked at the NS effective-surface range due to the function $\mathcal{D}_{1}$ [Eqs. (29)-(30)]. Therefore, one finds the equation of state in the form:
@@
Figure 6: The same as in Fig. 5 but for the total pressure $p(r,A_{\rm S})$, Eq. (50), as function of the radial coordinate $r$ and the dimensionless parameter $A_{\rm S}$ of the Schwarzschild metric, Eq. (II). The numbers in squares show again the values of this pressure. White color presents regions where we have zero values. The effective NS radius $R=10$ km is the same as in Figs. 1-4. The leptodermic parameter has a typical value $a/R=0.1$. For example, the parameter $\kappa\equiv\overline{\rho}K_{G}/12\mathcal{E}_{0}=10$ for large gravitational forces ($\kappa\gg 1$).
Figure 7: The dimensionless pressure $p=P(8\pi GR^{2}/c^{4})$, Eq. (50), as function of the radial coordinate $r$ for the Schwarzschild metric in the form of Eq. (II) at the same value of the parameter $A_{\rm S}$, $A_{\rm S}\approx 1.15$, and different values of the incompressibility parameter $\kappa=0$ [black solid for only zero order pressure $p_{0}$, Eq. (47)], $1$ [red dashed line, Eq. (50)], $5$ (green dotted line), and $10$ (blue dash-dotted line). The NS mass is the same as for the solid line in Fig. 3, $M=1.4M_{\odot}$. Other parameters are the same as in Fig. 6.
Figure 8: The same as in Fig. 7 but for the NS mass presented as the dashed line in Fig. 3, $M=2.0M_{\odot}$ ($A_{\rm S}=0.96$). Other parameters are the same as in Figs. 6 and 7.
@@
$K$ is a non-gravitative (e.g., nuclear) incompressibility; see also Eqs. (6) for $\mathcal{E}_{0}$ and (18) for a smooth part of the energy density $\mathcal{A}(\rho)$ in terms of the gravitational potential derivatives of $\Phi$. For the total dimensionless pressure $p$, at first order in the leptodermic parameter $a/R$, one finally obtains

### v2

@@ II Tolman-Oppenheimer-Volkoff approach
Notice that the solution, Eq. (11), coincides, up to a constant, with that presented in Tolman’s book at $\zeta=B_{\rm S}/A_{\rm S}=1/(2A_{\rm S})$ (see Refs. RT87 ; MM24 and Eq. (9) for $R_{\rm S}$). The cosmological constant also is assumed to be zero, and the units for the dimensionless pressure are used, $\mathcal{E}_{0}\propto R^{-2}_{\rm S}$, Eq. (9). Thus, one obtains the relation of these parameters of the Schwarzschild metric, Eq. (II), to the initial condition $P=P_{0}$ at $r=0$ through Eq. (12) for $\zeta$. Notice that according to Eq. (11) for the pressure $P$, one has $P=0$ at the boundary $r=R$ because of Eq. (10) for the constants $A_{\rm S}$ and $B_{\rm S}$. Finally, our result for the pressure, Eq. (11), coincides identically with that of Ref. RT87 .
Figure 2 shows the contour plots for the pressure $P(r)$ [in units of $P(r=0)$], as function of $r$ (km) and parameter $A_{\rm S}$, Eq. (10). It is more convenient to use the finite dimensionless values of $A_{\rm S}$, $0\leq A_{\rm S}\leq 1.5$, instead of dimensionless $R_{\rm S}/R$ for the fixed effective NS radius $R=10$ km, $0\leq r\leq R$. For a very rough estimate neglecting the relatively small gravitational mass defect, $\mathcal{E}_{0}\sim c^{2}\rho^{(0)}$, where the central mass density $\rho^{(0)}\approx M/(4\pi R^{3}/3)\sim M_{\odot}/(4\pi R^{3}/3)$, ($M\approx M_{V}$) and $M_{\odot}=1.989*10^{30}$ kg is the Sun mass, one finds significantly larger Schwarzschild radius parameter $R_{\rm S}$ than the effective radius $R$, $R_{\rm S}\approx 15.6$ km for $M=1.4M_{\odot}$, as example. By definition, $R_{\rm S}$ is always larger than (or equal to) the effective NS radius $R$, $R\leq R_{\rm S}$, while the gravitational radius $r_{\rm g}$ should be smaller than $R$, $r_{\rm g}=2GM/c^{2}\leq R$ LLv2 . These conditions are fulfilled in our calculations for the values $R_{\rm S}\approx 15.6$ km, and $r_{\rm g}\approx 4.1$ km in the case of the effective radius $R=10$ km. With these evaluations, one has the constant $A_{\rm S}\approx 1.15$. The maximum value $A_{\rm S}\approx 1.5$ corresponds to zero pressure $P(r)$.
@@ III Modified TOV approach
We now derive the leptodermic corrections to the TOV equations due to the gradient terms of the energy density $\mathcal{E}$, in line of Ref. MM24 . The details of these derivations are shown in Appendix A.
@@ III.1 General points
The key quantity of our derivations is the energy density $\mathcal{E}\left(\rho\right)$ strmagden ; BM15 ; MM24 ,
@@
The volume element, $\hbox{d}\mathcal{V}$, for a spherical system is determined by the Schwarzschild metric, $\hbox{d}\mathcal{V}=\hbox{exp}(\lambda/2)\hbox{d}{\bf r}$, where $\hbox{d}{\bf r}=r^{2}\hbox{sin}\theta~\hbox{d}\theta\hbox{d}\varphi$. The integration over $\hbox{d}\mathcal{V}$ is carried out for the whole infinite space. In Eq. (13), $\mathcal{A}(\rho)$ is the volume non-gradient part, Eq. (9),
@@
where $K$ is the non-gravitational part of the incompressibility modulus, e.g., due to the Skyrme nuclear interaction (for nuclear matter $K\sim 200$ MeV), $\Phi(\rho)$ is the statistically averaged (macroscopic) gravitational field; see Eq. (8) and discussions around it in Ref. MM24 . For simplicity, we will consider the gradient term of the inter-particle interaction with a main constant $\mathcal{C}$ of Eq. (12) for a smooth coefficient $\mathcal{B}(\rho)$ in front of the gradient-density square term.
@@
Therefore, when we can neglect external forces, there is no linear terms in expansion of $\mathcal{A}$, Eq. (15), over powers of the difference $\rho-\overline{\rho}$ near the saturation value $\overline{\rho}$. Near the saturation value, $\rho\rightarrow\overline{\rho}$, one has $\mathcal{W}=\mathcal{A}/\rho$. Expanding the gravitational field $\Phi(\rho)$ near the saturation density $\overline{\rho}$ up to second order, and taking into account Eq. (17), one indeed finds that there is no linear term, namely, $\partial\Phi/\partial\rho=0$ at $\rho=\overline{\rho}$, i.e., $\Phi=\Phi_{0}+(1/2)\Phi_{2}(\rho-\overline{\rho})^{2}$ with some constants $\Phi_{0}$ and $\Phi_{2}$. Combining different terms in Eq. (15), one writes
$\mathcal{A}=-b^{(G)}_{V}\rho+\varepsilon_{G}(\rho),$ (18)
where $b^{(G)}_{V}=b_{V}-m\Phi_{0}$, $m$ is the test particle mass, $\Phi_{0}=\Phi(\overline{\rho})$. The particle separation energy, $b^{(G)}_{V}$, and energy density component, $\varepsilon_{G}(\rho)$ (see Eq. (16) for $\varepsilon(\rho)$) are modified both by the gravitational field $\Phi(\rho)$,
@@
$\hskip-14.22636pt\frac{\delta\mathcal{E}}{\delta\rho}\equiv\frac{\partial\mathcal{A}}{\partial\rho}-2\mathcal{C}\Delta\rho=\mu~.$ (22)
@@
The solution of the Lagrange equation (22) at leading order of the leptodermic expansion over the parameter $a/R$, $\rho(r)=\overline{\rho}y((r-R)/a)$, as function of the radial coordinate $r$, is shown in Fig. 1 in the main case (i), $\mathcal{B}=\mathcal{C}=const$, for typical parameters, the thickness of the NS diffused layer (crust) $a=1$ km and the effective radius $R=10$ km, in units of the saturation value $\overline{\rho}$; see Refs. ST04 ; HPY07 ; MM24 . We compare the two dimensionless solutions for the particle density $\rho(r)/\overline{\rho}$. One of them (solid line) is related to the vdW-Skyrme interaction constant $\mathcal{C}$ “1” for a dense liquid drop; see Eq. (32) for the universal dimensionless particle number density of a dense liquid drop, $y=\rho/\overline{\rho}$, as function of x=(rR)/ax=(r-R)/a. Another limit case corresponds to the symmetric Wilets solution (ii) wilets ; MM24 for zero constant $\mathcal{C}$ and no spin-orbit term but with taking into account the only kinetic gradient term in a gas system; see Eq. (12) for $\mathcal{B}$. These two curves are very different in the surface layer, especially outside of the system far away from the NS, and almost the same inside of the NS. As seen from Fig. 1 and expected, the range of the exponential decrease of $\rho(r)$ becomes much smaller for the vdW-Skyrme liquid-drop case (i) than for the Wilets gas case (ii). But the surface effects on the local density $\rho(r)$ in both cases are still very important to get a continuous analytical behavior through the ES. The smallness factor, proportional to $a/R$, with respect to the volume contribution, appears because of the integration over the normal-to-ES (radial in this example) variable x=(rR)/ax=(r-R)/a in this ES region. Thus, for the calculations of the surface corrections to the NS total energy $E$, Eq. (14), and particle number $N$, Eq. (21), it would seem possible to assume that they are not (very) important for so small relative diffuseness thickness $a/R=0.1$. However, as shown in Ref. MM24 , the influence of the NS surface becomes really important for the local quantities as the particle density $\rho(r)$, Fig. 1, the EoS, and therefore, for the pressure $P(r)$ calculating through Eqs. (13) and (17),
@@
Moreover, the gravitational mass defect due to the radial curvature of the nonlinear Schwarzschild metric in calculations of the NS mass leads to its dramatic non-monotonic behavior of $M=M(R)$; see Ref. MM24 .
@@ III.2 Surface corrections to the TOV equations
Indeed, $\mathcal{E}_{1}$ enters the solutions of Eq. (II) in terms of the integral over the radial coordinate $r$. Our derivations at leading order over leptodermic parameter $a/R$ are similar to those for the equation for the density $\rho$ and its solutions in Ref. MM24 .
Starting from the same equations (II), as presented in the previous section (see Ref. RT87 ), we split all their solutions into two components, the leading component and leptodermic corrections,
@@
As shown in Appendix A, using approximately the property of the energy density component $\mathcal{E}_{1}\propto(\partial\rho/\partial r)^{2}$, which is concentrated near the ES, for the correction $\lambda_{1}$, at first order in the leptodermic expansion, one finally obtains from the second equation of Eq. (II),
@@
The integrations over $r$ and $\xi=r-R$ are indefinite, as well as in all equations below, in particular in Appendix A.
Substituting the asymptote of the solutions, $\rho\sim\hbox{exp}(-\xi/a)$, for larger distances $\xi/a\gg 1$ (see Ref. MM24 ), one has an exponential disappearance of the integral $\mathcal{D}_{1}$, Eq. (29),
@@
For the inner region, one finds also the exponentially decreasing asymptotes of $\mathcal{D}_{1}$, Eq. (29),
@@
where $y=\rho/\overline{\rho}$ (see Ref. MM24 ), $C_{y}$ is an integration constant, and $\hbox{d}y=\hbox{d}r(\hbox{d}y/\hbox{d}x)/a$ with $y=y(x)$ of Eq. (32) in the integral, Eq. (III.2). To remove singularity at the origin, we shall assign the value 4/15-4/15 for the constant $C_{y}$. Notice that in the derivations for $\lambda_{0}$, Eq.(27) [Eq. (96.7) in Tolman’s book RT87 ] similar corrections were neglected by the same arguments. The solution $\lambda_{1}(r)$, Eq. (28), is zero outside of the system because of zero $\mathcal{E}_{0}$, Eq. (6), and $P_{0}$, Eq. (11). Therefore, $\lambda_{1}(r)$ is concentrating near the NS surface, as it should be true (see also Appendix A).
Substituting the derivative ${\rm d}\nu/{\rm d}r$ from the first equation of Eq. (II) to its last equation, for a given energy density $\mathcal{E}$, one obtains the modified TOV equation for the pressure $P$,
@@
where $\lambda=\lambda_{0}+\lambda_{1}$; see Eqs. (27) and (28). Notice that for shortness, from this equation to the final result, Eq. (45), and in Appendix A we use the generally accepted units $c=G=1$, in particular as in Ref. RT87 . Equation (33) is valid approximately up to high, e.g. second order in the leptodermic expansion over $a/R$. Let us use now Eqs. (24) for $\mathcal{E}$ and (26) for $P$. Using also Eq. (A1) for $P_{0}$ at zero order of the leptodermic parameter $a/R$, one can derive the linearized equation for $P_{1}$ beyond the analytical solutions (27) and (11) of the standard TOV equations for $\lambda_{0}$ and $P_{0}$, respectively.
@@
where $\tilde{r}=r/R_{\rm S}$, as in Appendix A. In these equations, $P_{0}(r)$ is the analytical solution of the TOV equations at zero order of the leptodermic parameter for the pressure, Eq. (11); see Refs. RT87 ; MM24 . Using the same method of the arbitrary constant variations, one obtains the solution in terms of the indefinite integrals quadratures (see Appendix A),
@@ III.3 Leading solutions for the pressure
Substituting the function $\beta(r)$, Eq. (35), into the indefinite integrals for the pressure correction $P_{1}$, Eqs. (37) and (38), one finds
@@
where $\zeta=1/(3\sqrt{1-\tilde{R}^{2}})$ with $\tilde{R}=R/R_{\rm S}$. Using then Eq. (36) for $\beta_{1}(r)$ in Eq. (38) for $C_{1}^{(P)}$, one obtains the sum of the two terms,
@@
The integrand of one of such integrals is proportional to $\mathcal{E}_{1}$, Eq. (25), while the second one is proportional to $\lambda_{1}$, Eq. (28). The first order corrections, $\mathcal{E}_{1}$ and $\lambda_{1}$, are determined by Eqs. (25) and (28), respectively [see also the approximate Eq. (A6)]. Both these functions exponentially decrease for $r$ outside of a relatively small range $a$ of the NS crust at small leptodermic parameter, $a/R$. Therefore, the first integral in Eq. (III.3) can be taken approximately at leading order, $a/R$, in the leptodermic expansion as for the derivation of $\lambda_{1}$. Notice that the second integral in Eq. (III.3), with the integrand $\propto\lambda_{1}(r)$, over radial coordinate $r$ can be neglected as the second order of $(a/R)^{2}$ in the leptodermic expansion. Therefore, it does not contribute at first order over $a/R$. Indeed, this is reduced approximately to $\int{\rm d}r\mathcal{Y}(y)$,
@@
$\times\int{\rm d}r\mathcal{Y}(y).$ (41)
Notice that by the same reason, the $\lambda_{1}$ surface correction to the mass defect owing to the curved space can be neglected in calculations of the NS mass $M$ (Ref. MM24 ) at the first order over $a/R$. Transferring the last integral over $r$ to the variable $\rho$ and then to the dimensionless $y$, as above, ${\rm d}r=a{\rm d}y/({\rm d}y/{\rm d}x)$, and using Eqs. (27), and (19) for the inverse derivative of $y$ over $x$, one finds
@@
The indefinite integral on the right hand side is finite because its integrand has no singularity in the limit $y\rightarrow 1$ by the L’Hopital theorem, $\mathcal{Y}/(1-y)\rightarrow$ $\mathcal{Y}^{\prime}(y)/(-1)\rightarrow(1-y)\rightarrow 0$~. There is also no singularity at small $y$, $\int{\rm d}y/\sqrt{y}=2\sqrt{y}$.
@@
where $\mathcal{D}_{1}$ is given by Eq. (29) with the same asymptotic properties (30) and (31) outside of the narrow NS crust, and again, $\tilde{R}=R/R_{\rm S}$. With Eq. (39) for the integral exponent taken at $r=R$ and the coefficient $C^{(P)}_{1}$, Eq. (43), one arrives at the final analytical approximation for the surface pressure correction $P_{1}$,
$P_{1}=-\frac{3R^{5/2}}{4R_{\rm S}^{5/2}\sqrt{1-R^{2}/R^{2}_{\rm S}}}~\frac{\mathcal{D}_{1}}{R}~;$ (44)
see Eq. (37). This final result (and below) is presented in terms of the coefficient $\mathcal{D}_{1}/R$ which has the clear pressure dimension [see Eq. (29)]. Another factor is dimensionless as that expressed through the ratio $R/R_{\rm S}$ of the effective radius $R$ to the Schwarzschild radius $R_{\rm S}$, Eq. (9), due to the Schwarzschild gravitation. Therefore, taking Eq. (9) into account, one can easy come back from the $c=G=1$ to the normal units. The Schwartzschild radius $R_{\rm S}$, Eq. (9), is related to the initial value of the energy density, $\mathcal{E}_{0}$, and $\mathcal{D}_{1}$, Eq. (29), has the internal (31) and external (30) asymptotes. Therefore, these leptodermic corrections for $C_{1}^{(P)}$, and hence, for $P_{1}$, (as $\propto\mathcal{D}_{1}$), are exponentially decreasing functions of radial coordinate $r$, $r\rightarrow\infty$ outside [see Eq. (30)] and to zero inside of the system [Eqs. (31) and (III.2)]. In particular, inside of the system, Eq. (31) for $\mathcal{D}_{1}$, one arrives at
$P_{1}=\frac{a}{R}~\frac{\overline{\rho}K_{G}}{12}~\frac{R^{5/2}}{R^{5/2}_{\rm S}\sqrt{1-R^{2}/R^{2}_{\rm S}}}~\mathcal{Y}(y)~,$ (45)
with the explicitly presented pressure dimension. Thus, as expected, the leading leptodermic correction $P_{1}$ to the zero order pressure $P_{0}$ is peaked at the NS effective-surface range due to the function $\mathcal{D}_{1}$ [Eqs. (29)-(30)]. Therefore, one finds the equation of state in the form:
@@
Figure 6: The same as in Fig. 5 but for the total pressure $p(r,A_{\rm S})$, Eq. (50). White color presents regions where we have zero values. The effective NS radius $R=10$ km is the same as in Figs. 1-4. The leptodermic parameter has a typical value $a/R=0.1$. For example, the dimensionless incompressibility parameter $\kappa=10$ [see Eq. (49)] for large gravitational forces ($\kappa\gg 1$).
Figure 7: The dimensionless pressure $p=P(8\pi GR^{2}/c^{4})$, Eq. (50), as function of the radial coordinate $r$ for the Schwarzschild metric in the form of Eq. (II) at the same value of the parameter $A_{\rm S}$, $A_{\rm S}\approx 1.15$, and different values of the incompressibility parameter $\kappa=0$ [black solid for the zero order pressure $p_{0}$, Eq. (47)], 11 [red dashed line, 55 (green dotted line), and 1010 (blue dash-dotted line), Eq. (50)]. The NS mass is the same as for the solid line in Fig. 3, $M=1.4M_{\odot}$. Other parameters are the same as in Figs. 5 and 6.
Figure 8: The same as in Fig. 7 but for the NS mass presented by the dashed line in Fig. 3, $M=2.0M_{\odot}$ ($A_{\rm S}=0.96$).
@@
$K$ is a non-gravitating (e.g., nuclear) incompressibility; see also Eqs. (6) for $\mathcal{E}_{0}$ and (18) for a smooth part of the energy density $\mathcal{A}(\rho)$ in terms of the gravitational field derivatives of $\Phi$. For the total dimensionless pressure $p$, at first order in the leptodermic parameter $a/R$, one finally obtains